A block with a mass of 0.20 kg is hung at the end of a spring and it is lifted vertically so that the spring is unstretched. The block is then realesed and it touches the table with a velocity of 0.80 m s-1. If the distance of the table from the realesed point of the spring is 0.10 m, what is the force constant of the spring?


1
Expert's answer
2021-11-25T16:23:41-0500

"E_1=E_2,"

"\frac{kx^2}2=\frac{mv^2}2,"

"k=\frac{mv^2}{x^2}=1.28~\frac Nm ."